OpenStudy (anonymous):

The side length of a square is (x - 4)^2 units. What is the perimeter of the square? * x^2 - 8x + 16 4x^2 - 32x + 64 x^2 + 16 4x^2 + 64

OpenStudy (anonymous):

Wow harddd stuff

OpenStudy (anonymous):

whats the answer

OpenStudy (zzr0ck3r):

perimeter of a square with side \(x\) is \(4x\) so \(4(x-4)^2 = 4(x^2-8x+16) = \ \large ? \)

OpenStudy (anonymous):

whats the answer???